# Guaranteed confidence-band enclosures for PDE surrogates

### v1

@@ Acknowledgements
Appendix for “Guaranteed confidence-band enclosures for PDE surrogates”
@@ Appendix B Relationships between set-representations used in the paper
We find the LazySets.jl [^2] https://github.com/JuliaReach/LazySets.jl (Forets and Schilling, 2021) manual, and software docstrings, quite thorough resources for set-representations and their respective overapproximations. Here we summarise some set isomorphisms and overapproximations used the main paper.
@@ B.2 Overapproximating a polytope with a zonotope
With the algorithm initially proposed by guibas2003zonotopes (section 4.2), here we summarise the version implemented in LazySets.jl (Forets and Schilling, 2021). Further detail can be found therein.
@@
Appendix C Fourier Neural Operators, Burgers’ equation, and Training
Here we provide extra details about the example PDE, the functional surrogate model, and its training, used in the main paper.
C.1 Burgers’ Equation
@@
A dataset of $2048$ (training) + $1000$ (calibration) + $1048$ (validation) PDE solutions is generated by Latin Hypercube sampling (uniform) the $\alpha$, $\beta$, and $\gamma$ parameters (thus generating random initial conditions for the PDE), and then solving Burgers’ equation using a spectral solver (canuto2007spectral). Each simulation is run for $500$-time iterations with a $\Delta t=0.0025$ time step and a spatial domain spanning $[0,1]$, uniformly discretised into $1024$ spatial units. The field at the last time point is then saved as the output, and the surrogate’s task is to learn the mapping from the initial condition to the fields final state $u(y,0)\rightarrow u(y,t_{\text{end}})$.
C.2 Fourier Neural Operators
Fourier Neural Operators (FNOs), introduced by Li et al. (2020), are specific instance of the Neural Operator (NO) class of ML models, which have shown efficacy in mapping between function spaces. Following a description by Gopakumar et al. (2024), a NO can be written as a parameterised mapping between function spaces $G_{\theta}:A\mapsto U$, where $G_{\theta}$ is a neural network parameterised by $\theta$, with three specific architecture elements, which are sequential:
@@ C.2 Fourier Neural Operators
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ are the Fourier and inverse Fourier transform, and $R_{\phi}$ is a learnable complex-valued tensor comprising of truncated Fourier modes. In practical application, the network is discretised to a finite number of modes and the discrete FFT is used for $\mathcal{F}$. FNOs are learnable using gradient based optimisation using automatic differentiation.
C.3 Training

### v2

@@ Appendix B Relationships between set-representations used in the paper
We find the LazySets.jl [^1] https://github.com/JuliaReach/LazySets.jl [Forets and Schilling, 2021] manual, and software docstrings, quite thorough resources for set-representations and their respective overapproximations. Here we summarise some set isomorphisms and overapproximations used the main paper.
@@ B.2 Overapproximating a polytope with a zonotope
With the algorithm initially proposed by Guibas et al. [2003] (section 4.2), here we summarise the version implemented in LazySets.jl [Forets and Schilling, 2021]. Further detail can be found therein.
@@
Appendix C Extended experiment results
Table 2 gives further information about the data sets, models, and output dimensions for the experiments in this paper. We note that the MLP benchmarks (Bias to Crime) were originally found in Messoudi et al. [2022] for their comparison of multivariate conformal prediction. Table 3 gives an extended presentation of experimental results, including those for lower dimensional MLPs. Note that for the non-PDE benchmarks, the prediction sets’ volume was used as an efficiency metric, while for the high dimensional problem we used average 2D areas, as detailed in the main text.
Table 2: Information about multivariate and PDE datasets
Name model Calibration data dimensions Source
Bias MLP 7750 2 Cho et al. [2020]
Music MLP 350 2 Zhou et al. [2014]
Indoor MLP 6946 2 Torres-Sospedra et al. [2014]
SGMM MLP 79728 4 Nugteren and Codreanu [2015]
Crime MLP 628 18 Redmond and Baveja [2009]
Burgers FNO 2048 1024 Gopakumar et al. [2024a]
Burgers DeepONet 2048 1024 Gopakumar et al. [2024a]
Wave FNO 7000 4096 Gopakumar et al. [2024a]
Navier Stokes FNO 7000 4096 Gopakumar et al. [2024a]
Table 3: Results of coverage and efficiency comparison on multivariate and PDE data sets.
$\alpha=0.1$ Coverages Efficiency
$\alpha=0.2$ Modulation Copula Ellipse Rotated Zonotope Modulation Copula Ellipse Rotated Zonotope
dataset / PDE (model) Box (ours) (ours) Box (ours) (ours)
Bias $89.46$ $89.06$ $91.13$ $88.18$ $91.29$ $\mathbf{5.186^{-2}}$ $5.478^{-2}$ $5.418^{-2}$ $5.460^{-2}$ $6.280^{-2}$
$78.99$ $79.95$ $81.15$ $79.15$ $82.59$ $\mathbf{3.346^{-2}}$ $3.508^{-2}$ $3.357^{-2}$ $3.525^{-2}$ $3.913^{-2}$
Music $89.71$ $94.86$ $96.57$ $90.86$ $90.29$ $\mathbf{4.927^{-1}}$ $5.504^{-1}$ $7.467^{-1}$ $6.905^{-1}$ $6.125^{-1}$
$81.71$ $86.29$ $86.29$ $86.29$ $82.29$ $\mathbf{3.235^{-1}}$ $3.460^{-1}$ $4.104^{-1}$ $5.512^{-1}$ $4.425^{-1}$
Indoor $90.64$ $90.96$ $90.18$ $88.80$ $91.22$ $8.105^{-2}$ $8.655^{-2}$ $\mathbf{7.377^{-2}}$ $1.217^{-1}$ $1.377^{-1}$
$80.33$ $81.11$ $81.03$ $78.72$ $81.20$ $\mathbf{4.494^{-2}}$ $4.580^{-2}$ $4.523^{-2}$ $7.108^{-2}$ $7.473^{-2}$
SGMM $89.98$ $90.71$ $89.48$ $89.88$ $89.85$ $2.541^{-5}$ $2.950^{-5}$ $\mathbf{1.176^{-9}}$ $2.224^{-6}$ $2.632^{-6}$
$80.01$ $80.93$ $79.37$ $79.74$ $80.28$ $4.142^{-6}$ $4.868^{-6}$ $\mathbf{2.331^{-10}}$ $1.444^{-6}$ $1.715^{-6}$
Crime $92.04$ $89.17$ $90.45$ $87.58$ $91.08$ $4.125^{-10}$ $8.923^{-7}$ $\mathbf{1.583^{-22}}$ $6.705^{-18}$ $7.739^{-15}$
$85.35$ $79.30$ $80.89$ $78.66$ $80.89$ $1.387^{-12}$ $1.889^{-9}$ $\mathbf{4.171^{-25}}$ $8.472^{-19}$ $2.793^{-15}$
Burgers (FNO) $89.36$ $91.41$ $88.57$ $3.956$ $2.927^{-1}$ $\mathbf{1.910^{-1}}$
$78.32$ $83.98$ $79$ $1.634$ $2.562^{-1}$ $\mathbf{1.867^{-1}}$
Burgers (DeepO) $90.92$ $92.77$ $89.65$ $2.513^{3}$ $\mathbf{5.868}$ $8.330$
$80.96$ $83.01$ $79.69$ $9.899^{2}$ $\mathbf{4.052}$ $7.580$
Wave (FNO) $86$ $89.2$ $90$ $3.588^{4}$ $5.656^{-3}$ $\mathbf{5.258^{-3}}$
$78$ $78.4$ $81.6$ $3.559^{4}$ $\mathbf{4.569^{-3}}$ $4.744^{-3}$
Navier Stokes $86.83$ $89.17$ $87.33$ $9.251^{4}$ $\mathbf{3.423^{1}}$ $3.529^{1}$
(FNO) $75.83$ $78.67$ $79.67$ $7.040^{4}$ $\mathbf{2.441^{1}}$ $3.282^{1}$
$X^{Y}$ denotes $X\times 10^{Y}$ and "—" denotes no result
Appendix D Physics, Surrogate Models, and Training
Here we provide extra details about the PDEs, the functional surrogate models, and their training configurations, used in the main paper. The numerical solvers, data and functional surrogates models used for PDE modelling is borrowed from [Gopakumar et al., 2024a].
D.1 Burgers’ Equation
@@
A dataset of $2048$ (training) + $1000$ (calibration) + $1048$ (validation) PDE solutions is generated by Latin Hypercube sampling (uniform) the $\alpha$, $\beta$, and $\gamma$ parameters (thus generating random initial conditions for the PDE), and then solving Burgers’ equation using a spectral solver [Canuto, 2007]. Each simulation is run for $500$-time iterations with a $\Delta t=0.0025$ time step and a spatial domain spanning $[0,1]$, uniformly discretised into $1024$ spatial units. The field at the last time point is then saved as the output, and the surrogate’s task is to learn the mapping from the initial condition to the fields final state $u(y,0)\rightarrow u(y,t_{\text{end}})$.
D.2 Wave Equation
$\frac{\partial^{2}u}{\partial t^{2}}=c^{2}\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\right),$
where $u$ defines the field variable, $c$ the wave velocity, with $x$, $y$ and $t$ being the spatial and temporal coordinates respectively. The initial conditions are defined as:
$u(x,y,t=0)=\exp(-\alpha((x-\beta)^{2}+(y-\gamma)^{2})),$
parameterised by $\alpha\in[10,50]$, $\beta\in[0.1,0.5]$, and $\gamma\in[0.1,0.5]$, with an additional constraint $\frac{\partial u}{\partial t}(x,y,t=0)=0$.
Data set generation
A dataset of $500$ (training) + $1000$ (calibration) + $1000$ (validation) PDE solutions is generated by Latin Hypercube sampling the $\alpha$, $\beta$, and $\gamma$ parameters. The wave equation is solved using a spectral solver with leapfrog time discretization and Chebyshev spectral method on tensor product grid [Gopakumar et al., 2023]. Each simulation runs for $150$ time iterations with $\Delta t=0.00667$ across a spatial domain of $[-1,1]^{2}$, discretized into $33$ spatial units per dimension. The first $80$ time instances of each simulation are used for training.
D.3 Navier-Stokes Equations
The Navier-Stokes scenario that we are interested in modelling is taken from the exact formulation in Li et al. [2020], where the viscosity of the incompressible fluid in 2D is expressed as:
$\partialderivative{w}{t}+u\nabla w$ $=\nu\nabla^{2}w+f,$ $\quad x\in(0,1),\;y\in(0,1),\;t\in(0,T)$ (5)
$\nabla u$ $=0,$ $\quad x\in(0,1),\;y\in(0,1),\;t\in(0,T)$ (6)
$w$ $=w_{0},$ $\quad x\in(0,1),\;y\in(0,1),\;t=0,$ (7)
where $u$ is the velocity field and vorticity is the curl of the velocity field $w=\nabla\crossproduct u$. The domain is split across the spatial domain characterised by $x,y$ and the temporal domain $t$. The initial vorticity is given by the field $w_{0}$. The forcing function is given by $f$ and is a function of the spatial domain in $x,y$. We utilise two datasets from Li et al. [2020] that are built by solving the above equations with viscosities $\nu=1e-3$ and $\nu=1e-4$ under different initial vorticity distributions. For further information on the physics and the data generation, refer Li et al. [2020].
D.4 Multilayer Perceptrons
Multilayer Perceptrons (MLPs), a fundamental class of neural networks, are composed of sequential layers of neurons that transform input features through learned weight matrices and non-linear activations [Haykin, 1994]. For a given input $x\in\mathbb{R}^{d_{\text{in}}}$, an MLP with $L$ layers computes:
$h_{i+1}$ $=\sigma(W_{i}h_{i}+b_{i}),\quad i=0,\ldots,L-1$
$h_{L}$ $=W_{L}h_{L}+b_{L}$ (8)
where $h_{0}=x$, $W_{i}\in\mathbb{R}^{d_{i+1}\times d_{i}}$ and $b_{i}\in\mathbb{R}^{d_{i+1}}$ are learnable parameters, and $\sigma$ is a non-linear activation function (in this case, hyperbolic tangent).
D.5 Fourier Neural Operators
Fourier Neural Operators (FNOs), introduced by Li et al. [2020], are specific instance of the Neural Operator (NO) class of ML models, which have shown efficacy in mapping between function spaces. Following a description by Gopakumar et al. [2024b], a NO can be written as a parameterised mapping between function spaces $G_{\theta}:A\mapsto U$, where $G_{\theta}$ is a neural network parameterised by $\theta$, with three specific architecture elements, which are sequential:
@@ D.5 Fourier Neural Operators
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ are the Fourier and inverse Fourier transform, and $R_{\phi}$ is a learnable complex-valued tensor comprising of truncated Fourier modes. In practical application, the network is discretised to a finite number of modes and the discrete FFT is used for $\mathcal{F}$. FNOs are learnable using gradient-based optimisation using automatic differentiation.
D.6 MLP Training
The trained MLP model consists of:
  1. 1.
    Input layer: $\mathbb{R}^{d_{\text{in}}}\rightarrow\mathbb{R}^{256}$
  2. 2.
    Hidden layers: 3 layers of $\mathbb{R}^{256}\rightarrow\mathbb{R}^{256}$ transformations with tanh activation
  3. 3.
    Output layer: $\mathbb{R}^{256}\rightarrow\mathbb{R}^{d_{\text{out}}}$
The network was trained using the Adam optimizer with an initial learning rate of $5\times 10^{-3}$, which was reduced by a factor of 0.5 every 50 epochs. Training proceeded for 500 epochs using mini-batches of size 50 and mean squared error loss. Input features were normalized using a min-max scaling strategy applied per variable. The total parameter count for a single input/output dimension is $256(d_{\text{in}}+1)+256^{2}(L-1)+256+d_{\text{out}}(256+1)$.
D.7 FNO Training